OpenStudy (anonymous):

An archer releases an arrow from a shoulder height of 1.39 m. When the arrow hits the target 18 m away, it hits point A. When the target is removed, the arrow lands 45 m away. Find the maximum height of the arrow along its parabolic path. The distance from the top edge of the archery target to the ground is 1.5 meters. someone help??

OpenStudy (anonymous):

The target is 4 cm 4 cm 4 cm 4 cm and Point A on the target is 8 cm

OpenStudy (isaiah.feynman):

Strange wording.

OpenStudy (anonymous):

this question has been asked on here but nobody can clearly explain how its figured out im frustrated

OpenStudy (anonymous):

anybody know quadratic functions?

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

could you help me figure this out? want me to show what i have so far?

OpenStudy (anonymous):

y=a(0)^2+b(0)+c 1.39=c 1.5m-0.25m=1.26m 1.26=a(18)^2+b(18)+1.39 0=a(45)^2+b(45)+1.39 so the points (0,1.39);(18,1.26);(45,0)

OpenStudy (isaiah.feynman):

Sorry, what are those for?

OpenStudy (anonymous):

1.26m= Point A where the arrow hit and im trying to find the maximum height which from other threads ive seen the answer is 1.41 but doesnt really show how they got it ill post a picture of the actual question

OpenStudy (isaiah.feynman):

Okay the maximum height of the projectile WHEN the target is removed???

OpenStudy (anonymous):

It doesn't matter i just need the maximum height from when it was shot at 1.39m to when it hit point A The bulls eye on the target. i know its a confusing question

OpenStudy (anonymous):

But this was an added note not on the picture to help figure it out which was this: The distance from the top edge of the archery target to the ground is 1.5 meters. and if you subtract the 24 cm (0.24m) you get 1.26m

OpenStudy (isaiah.feynman):

Ok well its good to draw a diagram.I'm guessing its like this So the curved line will be the path of the arrow (which is a parabola) and we have to figure out its equation.